# SciNLP: A Domain-Specific Benchmark for Full-Text Scientific Entity and Relation Extraction in NLP

### v2

@@ 3.2 Comparison with Previous Datasets
Table 2: Examples of the relations types
@@ 4.2 Evaluation Settings and Results
Methods NER Rel Rel+ RE
Qwen3-8B(Pipeline) 58.56 18.76 16.76 21.56
Qwen3-8B(Joint) 52.63 19.62 17.18 -
DeepSeek R1-8B(Pipeline) 57.62 18.62 17.20 22.81
DeepSeek R1-8B(Joint) 51.42 14.62 11.38 -
Llama3.1-8B(Pipeline) 61.15 19.25 18.35 26.57
Llama3.1-8B(Joint) 53.15 14.26 11.15 -
Table 5: F1 scores of different LLM baselines on our proposed dataset.
@@ 5 Experiments
In this section, we utilize SciNLP to train SOTA supervised methods for SciNER and SciRE tasks, and compare them with existing datasets to validate the efficacy of our dataset. Section  5.1 introduces the supervised Baselines we have used. And Section  5.2 introduces the LLMs-based Baselines we have used. Section  5.3 presents the results of our dataset across these methods, followed by detailed discussions and analyses.
@@ 5.1 Supervised Baselines
PURE: (Zhong and Chen, 2021b) This model employs a dual-encoder architecture to decouple entity recognition and relation extraction tasks. The entity model performs span-based type prediction, while the relation model enhances classification through dynamic entity pair inputs and the insertion of entity-type markers.
PL-Marker: (Ye et al., 2022) The method proposes a floating tag mechanism: in the entity recognition stage, the tag and text span are dynamically bound by shared location coding and boundary awareness attention, and boundary learning is optimized by neighborhood clustering; In the relationship extraction stage, the subject-centered reorganization structure is used to jointly encode the semantics of subject and object, which overcomes the limitation of independent span to modeling. This unified method realizes the seamless integration of entity and relational representation.
HGERE: (Yan et al., 2023) Replacing strict NER modules with a lightweight pruning module, this method generates high-recall candidate entities, transferring classification tasks to a joint module to mitigate error propagation. It constructs hypergraph structures (comprising entity nodes, relation nodes, and multi-dimensional hyperedges) to aggregate high-order semantic information through hierarchical message passing.
5.2 LLMs-based Baselines
The adaptability of LLMs to diverse tasks enables them to efficiently handle a wide range of applications. This paper aims to evaluate the performance of LLMs on the SciIE task through carefully designed experiments. Leveraging the in-context learning (ICL) capabilities of LLMs, we designed both zero-shot and few-shot settings for pipelined extraction and joint ERE. Previous studies have shown that dynamically selecting few-shot demonstration examples for each test instance, rather than using a fixed set of contextual examples, can significantly enhance the performance of LLMs. Therefore, in our experiments, we employed a hybrid retrieval method to identify the most similar examples from the training set as in-context demonstrations for each given test sample.
For information extraction tasks, a well-designed prompt template can further improve the extraction performance of LLMs. Our prompt template primarily consists of the following components:
Task Description: The task instruction provides a basic description of the required task and specifies the output format that the model should follow.
Annotation Guidelines: These include standardized labeling instructions tailored for specific tasks such as NER, RE, and ERE, extracted from comprehensive annotation specifications. The guidelines cover, but are not limited to, entity definitions, relation definitions, etc., aiming to enable the LLM to perform annotation tasks accurately in accordance with the template instructions, similar to a human annotator.
Demonstrations: By computing sentence similarity scores using a retriever, relevant examples are selected from the training set to assist the model in better understanding the task requirements. Specifically, we select the top-k most similar examples ($x_{i}$, $y_{i}$) based on similarity scores to form the demonstration set $D$.
Test Case $x_{test}$: The input format is consistent with the examples $x_{i}$. We provide the input $x_{test}$ in the prompt template and guide the LLM to generate the corresponding output $y_{test}$ in the same format as the demonstrations.
Drawing on the experimental design techniques from SCIER, we also format NER annotations using HTML-style span tags. A complete example of the prompt template used in our experiments is provided in Appendix B. For the final experiments, we selected deepseek-r1-8b, llama3.1-8b, and qwen3-8b as the baseline LLMs for evaluation.
@@ 5.3 Result Analysis
Table  5 presents the experimental results of large language models (LLMs) in the few-shot setting. It can be observed that Llama3.1-8B outperforms Qwen3-8B and DeepSeek-R1-8B in overall performance. Furthermore, LLMs generally achieve better performance in pipelined information extraction tasks compared to Joint ERE methods. Specifically, in the NER task, the pipelined approach significantly outperforms the joint extraction method. This phenomenon suggests that splitting the end-to-end relation extraction task into separate NER and RE stages can enhance the overall performance of LLMs in this task.
@@ 6 Building Fine-grained KG in NLP Domain
Table  6 presents the statistics of nodes and relations in the KG we constructed. It can be observed that in the distribution of entity type quantities, entities of the "model" type account for a relatively high proportion. This indicates that NLP domain papers demonstrate a significant model-centric research paradigm, reflecting the core characteristics of rapid technological iteration and model diversification in this field. Additionally, the predominant occurrence of "usedFor" and "measuredBy" relationships highlights the characteristic of method and model-driven technological development in the NLP domain.
@@ A.3.1 Entity Relation Types
And for the relations, we have defined the following 11 entity relationships, Table  8 shows examples of correct relation annotations:

### v3

@@ 3.2 Comparison with Previous Datasets
Table 2: Examples of the relations types, the four entity types: Method, Metric, Dataset, Task.
@@ 5 Experiments
In this section, we utilize SciNLP to train SOTA supervised methods for SciNER and SciRE tasks, and compare them with existing datasets to validate the efficacy of our dataset. Section  5.1 introduces the supervised Baselines we have used. Section  5.2 presents the results of our dataset across these methods, followed by detailed discussions and analyses.
@@ 5.1 Supervised Baselines
PURE: (Zhong and Chen, 2021b) This model employs a dual-encoder architecture to decouple entity recognition and relation extraction tasks. The entity model performs span-based type prediction, while the relation model enhances classification through dynamic entity pair inputs and the insertion of entity-type markers. It introduces crosssentence context expansion by extending input windows, leveraging pre-trained models’ ability to capture long-range dependencies for improved global semantic reasoning.
PL-Marker: (Ye et al., 2022) Proposing a floating marker mechanism, this approach dynamically binds markers to text spans during the entity recognition phase through shared positional encoding and boundary-aware attention alignment, with neighborhood clustering further optimizing boundary learning. In the relation extraction phase, it restructures spans and markers via subject-centric reorganization, jointly encoding subject-object semantics to overcome limitations of independent span-pair modeling. This unified framework enables seamless integration of entity and relation representations.
HGERE: (Yan et al., 2023) Replacing strict NER modules with a lightweight pruning module, this method generates high-recall candidate entities, transferring classification tasks to a joint module to mitigate error propagation. It constructs hypergraph structures (comprising entity nodes, relation nodes, and multi-dimensional hyperedges) to aggregate high-order semantic information through hierarchical message passing. This design significantly outperforms traditional GNNs limited to pairwise interactions, particularly in capturing complex scientific logic across long documents.
All methods were implemented using the SciBERT-scivocab-uncased pre-trained model as the text encoder. For fair comparision, experiments adopted identical random seed sets to ensure robustness, along with consistent fine-tuning settings: learning rate = 2e-5, weight decay = 0.01, batch size = 16, and context length = 100.
@@ 6 Building Fine-grained KG in NLP Domain
Table  5 presents the statistics of nodes and relations in the KG we constructed. It can be observed that in the distribution of entity type quantities, entities of the "model" type account for a relatively high proportion. This indicates that NLP domain papers demonstrate a significant model-centric research paradigm, reflecting the core characteristics of rapid technological iteration and model diversification in this field. Additionally, the predominant occurrence of "usedFor" and "measuredBy" relationships highlights the characteristic of method and model-driven technological development in the NLP domain.
@@ 7 Conclusion
Ethical Statement
Our SciNLP dataset is constructed based on resources from The ACL OCL Corpus [^5] footnotetext: https://github.com/shauryr/ACL-anthology-corpus, a large corpus which provides full-text and metadata to the ACL anthology collection, and it is released under the CC BY-NC 4.0666footnotetext: https://creativecommons.org/licenses/by-nc/4.0/deed.en. All the other data are public from scientific documents. We release dataset for scientific information extraction tasks. There are no risks in our work.
In the process of the dataset construction, all human annotators are from our research group IR&TM, and all members participate voluntarily. The SciNLP dataset includes metadata extracted from the papers. No sensitive personal data (e.g., contact details or affiliations) is included. All metadata was collected in compliance with the terms of their sources and is used strictly for noncommercial academic research. We are committed to handling the data responsibly and ethically and will release our dataset under the same non-commercial license to ensure transparency and responsible data usage.
@@ A.3.1 Entity Relation Types
And for the relations, we have defined the following 11 entity relationships, Table  7 shows examples of correct relation annotations: